### v1

@@ 4 Conclusion
Chemical timescale reductions had negligible effect on large-scale macroscopic (peak pressure and wave speed) convergence characteristics for the cases considered here. Convergence of small-scale quantities was then assessed using instantaneous detonation profiles and relative locations of reaction zones in the wave structure. Supplementing the macroscopic convergence trends, the wave structure analysis revealed three characteristic detonation convergence regimes, namely the under-resolved, strongly-coupled, and ZND-conforming regimes. Convergence of reaction front locations uncovered the crucial role played by not only the chemical timescale, but also the ratio of chemical to induction timescales. The simplified cases exhibiting drastic reductions to this ratio showed not only delayed convergence to the strongly-coupled regime, but also significantly more deviation from baseline ZND solutions. The practical implication here is that if the induction timescale can be maintained, a reduction in chemical timescales due to mechanism simplifications and re-designs via combustion modeling can result in vastly improved detonation convergence properties.
Ultimately, this work highlights not only the important role played by the chemical timescale in detonation convergence, but also reveals the utility in assessing detonation convergence from the perspective of reaction front locations. The trends uncovered here open several avenues for extended verification, including performing the same analysis for different fuels, exploring how higher spatial dimensions modify the convergence regimes, and repeating the analysis with coupled time integration schemes that attempt to remove splitting errors. Further, this framework provides a methodology for mechanism reduction. Since shock-capturing requires somewhat stringent requirements as compared to subsonic flames, mechanism reduction could now focus on stiffness reduction within this constraint. In fact, related work suggests that machine learning models can be trained to enable precisely such approaches [33].
@@ Acknowledgments
The authors acknowledge funding through ONR MURI N00014-22-1-2606, with Dr. Steven Martens as program manager. SB acknowledges support from the AETS fellowship at Argonne National Laboratory under DOE contract DE-AC02-06CH11357.

### v2

@@ 4 Conclusion
Chemical timescale reductions had negligible effect on large-scale macroscopic (peak pressure and wave speed) convergence characteristics for the cases considered here. Convergence of small-scale quantities was then assessed using instantaneous detonation profiles and relative locations of reaction zones in the wave structure. Supplementing the macroscopic convergence trends, the wave structure analysis revealed three characteristic detonation convergence regimes, namely the under-resolved, strongly-coupled, and ZND-conforming regimes. Convergence of reaction front locations uncovered the crucial role played by not only the chemical timescale, but also the ratio of chemical timescale to induction time. The simplified cases exhibiting drastic reductions to this ratio showed not only delayed convergence to the strongly-coupled regime, but also significantly more deviation from baseline ZND solutions. The practical implication here is that if the induction time can be maintained, a reduction in chemical timescales due to mechanism simplifications and re-designs via combustion modeling can result in vastly improved detonation convergence properties. Ultimately, this work highlights not only the important role played by the chemical timescale in detonation convergence, but also reveals the utility in assessing detonation convergence from the perspective of reaction front locations.
The trends uncovered here open several avenues for extended verification and future work. A natural extension to the current study is to perform the same analysis, but for different fuels and mechanism settings. For example, investigating similar trends in convergence properties under chaotic and pulsating detonation simulation regimes (such as those uncovered in Ref. [58]) is a promising direction. Additionally, a direct way to build on the current study is to explore how higher spatial dimensions modify the convergence regimes, and the extent to which chemical timescale modifications can affect the cellular detonation structure. Considering that different cellular patterns may be obtained for different initial perturbations of the detonation wave, convergence of cell sizes may only be obtained in a statistical sense from an ensemble of simulations at the same resolution [59]. Even with the computational acceleration enabled by AMR, such an analysis would likely be extremely computationally expensive, especially when done in conjunction with investigations of chemical stiffness effects. In lieu of this, repeating the one-dimensional analysis in this study with coupled time integration schemes—instead of operator-split schemes—that attempt to remove splitting errors would lead to useful insights concerning the role played by chemical stiffness on the nature of detonation convergence. Such an analysis can build on previous work that has explored the relationship splitting errors and numerical solution deviations on different combustion problems [60].
Further, the framework in this study provides a methodology for mechanism reduction. Since shock-capturing requires somewhat stringent requirements as compared to subsonic flames, mechanism reduction could now focus on stiffness reduction within this constraint. In fact, related work suggests that machine learning models can be trained to enable precisely such approaches [61].
@@ Acknowledgments
The authors acknowledge funding through ONR MURI N00014-22-1-2606, with Dr. Steven Martens as program manager. SB acknowledges support from Argonne National Laboratory under DOE contract DE-AC02-06CH11357.